Shift Summary for evening shift, 23/06/99

Jianbo Wang (jianbo@fnal.gov)
Thu, 24 Jun 1999 00:10:33 -0500 (CDT)

Shift Summary for evening shift, 23/06/99

On Shift: Leo, Jianbo, Sydney

Hours of beam delivered at all intensities: 7.2
Hours of beam delivered at nominal intensity: 7.2
Hours of E832 data (special runs not included): 7.2
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

========================================================================
Runs ended this shift:
Run 13705 - 7.4M events - BEAM_832
Run 1370

Tapes brought to FCC this shift:
QKC 34 35 36
QKS 34 35 36
QKK 34 35 36
QKB 12

========================================================================
Accesses during shift:
No

========================================================================
Overview of shift:
E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:41 Stop run 13704, start 13705 in order to pick up
new filt832 namelist with relaxed T3 parameters
to prevent array over flows which induce missed
tracks.~

15:55 Beam is back in spill 4, BOA is 2.01

13705 16:00 Leo found the events.vs.spill is not flat. Call
MCR. After severals, the plots turn to be flat.

13705 167 19:42 cosmic#hodhits check found uneven hits. Leo found
that the last ped update done on June 19. We need
Ped run.

19:53 run 13705 stops baecause the tapes are full.

13706 20:05 Start run 13706

20:16 ktevana error, the ordinal number of HCC is not
correct, should be 9892, but it is 1892. run
kfrend, the problem is gone.

We still got lots of HCC inecciciency from chan
248. Val said it's not real inecciciency. We need
do some offline study.

13706 58 21:35 The BOA is 1.31.

21:49 Ktevana, run KFREND. Missing FFFF in crate 53 (in
fact, the order of stram 5 is screwed up in crate
53.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Please DON'T do muon run tonight. We lost a lot of data today, and there
will be shutdown tomorrow.